# On the Low-Temperature MCMC threshold:

### v2

@@ 4.3. The negative result
[Main negative result for sparse regression] Assume that for some sufficiently small $c>0$, we have
@@
Then w.h.p. as $k\rightarrow\infty$, for any reversible Markov chain on the graph $G_{p,k}$ with stationary distribution $\pi_{\beta}$, the mixing time of the chain is $\exp\left(\Omega(k\log(p/k))\right)$.
@@
To prove Theorem 4.4 (see Section 10), we establish a bottleneck for the MCMC methods, by significantly generalizing an Overlap Gap Property result from [GZ22a], where the authors assume $n=\Omega(k\log k)$ and $k=p^{o(1)}$. In particular, we generalize to arbitrarily small $n$ and any $k=o(p^{1/3})$.
@@ 5.2.1. Specification
Our implementation here uses $p=20000$, $k=18$, $\sigma^{2}=0.05$, and $\beta=100$. We then simulate the Metropolis chain for varying values of the sample size $n$. The information-theoretic threshold here is given by [GZ22a]
$n_{STATS}=\frac{2k\log(p/k)}{\log\left(\frac{k}{\sigma^{2}}+1\right)}\approx 43$
Our positive result in Theorem 4.2 provides the correct order for polynomial time convergence, but not the exact constant. We therefore compare to the algorithmic threshold for LASSO, where the exact threshold for success is given by [Wai09]
$n_{LASSO}=(2k+\sigma^{2})\log(p/k)\approx 253$
@@ 5.2.2. Results
The results of these simulations are displayed in Fig. 4. We again observe a clear phase transition when the sample size $n$ crosses the algorithmic threshold. When $n<n_{LASSO}$, most samples time out after a million iterations having failed to find $v^{*}$. Above $n_{LASSO}$, the chain quickly finds $v^{*}$. We can see these dynamics in more detail in Fig. 5. Here we group simulations by their value of $n$. Then at each time, we plot the median proportion of $v^{*}$ recovered at that time within that group. We observe that for values of $n$ above 300, we achieve very rapid convergence. Conversely, below 175 the chain does not appear to be converging at all, despite being significantly above the information-theoretic threshold $n_{STATS}\approx 43$.
We recognise that the correct constant for the MCMC algorithmic threshold may not be the same as for LASSO, and when $n$ is relatively small as it is here, that can affect the results. Nevertheless, we do see a rapid speed up in convergence when the LASSO threshold is surpassed.
@@
6. Further Comparison with Previous Work
6.1. Comparison with [AWZ23]
For sparse tensor PCA, our results directly compare with [AWZ23]. [AWZ23] studies the same setting exactly but only under the assumption $t=2$, i.e., in the matrix case, while we study the more general case $t\geq 2$. In this matrix case, they study the same class of low-temperature reversible MCMC methods on the Johnson graph as we consider in this work. They prove that whenever $k=o(p)$ their mixing time is super-polynomial for a part of the conjectured “computationally hard” regime that is when
The authors provide this as evidence for the conjectured computational hardness of sparse (matrix) PCA. Importantly they leave as an open question whether one of these MCMC methods mix in polynomial-time when other polynomial-time methods are known to work i.e. when $\lambda=\omega(\lambda_{\mathrm{ALG}}(2,k))$.
Our negative results prove that when $k\leq p^{\alpha_{0}}$ for $\alpha_{0}>0$ a small enough, a much wider failure takes place even when $t=2$. Specifically, all considered low-temperature MCMC methods in [AWZ23] have super-polynomial mixing time whenever $\lambda=o(k^{3/2}/\sqrt{\log p})$. Moreover, we also provide a positive result saying that mixing becomes polynomial for an element of this class (i.e., the Metropolis chain) when $\lambda=\omega(k^{3/2}\sqrt{\log p}).$ Hence we pin down the exact threshold for these MCMC methods to succeed in polynomial-time in sparse PCA, and we highlight that it turns out to be much bigger than $\lambda_{\mathrm{ALG}}(2,k)$.
From a technical standpoint, we obtain the mentioned stronger negative MCMC results compared to [AWZ23] because of a significantly improved application of the second moment method (see the proof of Proposition 9.4). The key ingredient we leverage is some novel “flatness” ideas from random graph theory [BBSV19, GZ19] appropriately adapted to Gaussian tensors. We expect this idea to be broadly useful for similar sparse problems.
6.2. Comparison with [GZ22a]
For sparse regression, the most relevant work is [GZ22a] that studies the exact same setting as we do. The authors prove that for all $k=o(p),$ if $n=\omega(n_{\mathrm{ALG}})$ a greedy local-search algorithm maximizing the posterior succeeds in polynomial-time. Moreover, they show that if $n=o(n_{\mathrm{ALG}})$ a disconnectivity property called Overlap Gap Property for inference takes place, which in particular leads to the failure of the mentioned greedy local search procedure. Importantly, their negative result applies only in the more restricted sparsity regime $k=p^{o(1)}.$
In our work, we prove more general both positive and negative results. For the positive side, we prove that for all $k=o(p),$ a larger family of low temperature MCMC methods succeeds whenever $n=\omega(n_{\mathrm{ALG}})$. In particular, our positive results include the randomized greedy algorithm maximizing the posterior, as opposed to the vanilla greedy approach of [GZ22a]. We highlight that this is a non-trivial extension, which is not always known to be true (see e.g., the relevant discussion in [GJX23] for the planted clique model) Finally, we prove as our negative result that for all $k=o(p^{1/3})$ a large family of low temperature MCMC methods fails when $n=o(n_{\mathrm{ALG}})$, significantly again generalizing the result of [GZ22a].
@@ 7. Proof of General Positive Result
For part (2), we use the standard canonical paths argument for mixing (for example, see chapter 5 of [Jer03]). For any $x,y\in\mathcal{X}$ we define $\gamma_{x,y}$ to be the unique path connecting them in $T$. Note that $|\gamma_{x,y}|\leq D$. Now for each edge $(a,b)$ in $T$, we upper bound the conductance $t(a,b)$ of it, defined as
@@
Thus, the conductance of the Metropolis chain on $G$ is at most $O(D\Delta)$ and the mixing time then follows, see [DS91, Sin92] or [Jer03, Theorem 5.2].
Lastly, we consider the case where $\beta\geq\frac{2}{\delta}(\log|\mathcal{X}|+\log(\frac{40\mathcal{R}_{H}}{\delta}))$ and prove part (3). We argue that in such a low-temperature regime, the behavior of the Metropolis chain is very similar as the Randomized Greedy algorithm with high probability. We call a step of the Metropolis chain to be “good” if, at current state $X_{t-1}=x\neq v^{*}$, it selects $y\in\Gamma_{G}(x)$ which is the parent of $x$ in $T$; note that since $H(y)>H(x)$ by 2.2, the chain will then move to $y$, i.e., $X_{t}=y$. For completeness, if the current state is $X_{t-1}=v^{*}$ then we always call this step good no matter which next state it picks or whether it moves. Thus, each step has a probability at least $1/\Delta$ to be good. A simple application of the Chernoff bounds yields that, if we run the Metropolis chain for $T=\frac{20\Delta\mathcal{R}_{H}}{\delta}$ steps, then at least $\frac{4\mathcal{R}_{H}}{\delta}$ of them are good with probability at least $0.9$.
Let $m^{\prime}$ be the number of good steps. Suppose the first $m=\min\{\frac{4\mathcal{R}_{H}}{\delta},m^{\prime}\}$ good steps happen at time $1\leq t_{1}<t_{2}<\cdots<t_{m}\leq T$; we also set $t_{0}=0$. Since the transition matrix, conditional on no move between time $t_{i-1}$ and $t_{i}$ is good, is still time-reversible, it holds that
@@
Now, by the union bound, with probability at least $0.8$, we have that (1) there are at least $m^{\prime}\geq\frac{4\mathcal{R}_{H}}{\delta}$ good steps; (2) for all $1\leq i\leq m=\frac{4\mathcal{R}_{H}}{\delta}$ it holds $H(X_{t_{i}-1})>H(X_{t_{i-1}})-\frac{\delta}{2}$. We claim that under (1) and (2), it must hold $X_{t_{i}-1}=v^{*}$ for some $1\leq i\leq m$, and the theorem then follows. Suppose for sake of contradiction that $X_{t_{i}-1}\neq v^{*}$ for all $i$. Then, by 2.2 we have $H(X_{t_{i}})\geq H(X_{t_{i}-1})+\delta>H(X_{t_{i-1}})+\frac{\delta}{2}$. It follows that $H(X_{t_{m}})>H(X_{t_{0}})+\frac{\delta m}{2}=H(X_{0})+2\mathcal{R}_{H}$, contradicting to $H(X_{t_{m}})-H(X_{0})\leq\mathcal{R}_{H}$ by definition. ∎
@@ 8.1. Proof of Positive Results for the Gaussian Additive Model
Here we apply Theorem 2.3 to prove Theorem 3.3 under Assumption 3.2. The first step is to show that Assumption 3.2 implies Assumption 2.2.
8.1.1. Verifying Assumption 2.2.
@@
In the following lemma, we verify Assumption 2.2 under Assumption 3.2.
@@
Assume that Assumption 3.2 holds with constants $\varepsilon\in(0,1)$ and $t\geq 1$. For all positive integers $p,k,\ell_{\min}\in\mathbb{N}^{+}$ such that $2/\varepsilon\leq k\leq p$ and $\ell_{\min}\leq\varepsilon k$, if
@@
We deduce from Assumption 3.2 that
@@
We want to show that there exists at least half of $y\in\Gamma_{G,v^{*}}(x)$ that satisfies $|Z^{x}_{y}|=O\left(\sqrt{(\log p)/k}\right)$. Observe that each $Z^{x}_{y}$ is a Gaussian with mean zero and variance given by
@@
where we use the convexity of $f$ and item (b) of Assumption 3.2. Thus, we have
@@
where we apply the convexity of $f$ and item (c) of Assumption 3.2. Meanwhile, the number of $y^{\prime}\in\Gamma_{G,v^{*}}(x)$ such that $\left\langle y,y^{\prime}\right\rangle=k-1$ is $2(k-\ell-1)$, and for such $w$ it holds
@@
by the convexity of $f$ and item (b) of Assumption 3.2. Finally, if $y^{\prime}=y$ then $\left\langle y,y^{\prime}\right\rangle=k$ and
@@
Lemma 8.2 (Hanson–Wright Inequality [RV13, Theorem 1.1]).
@@
We are now ready to prove Theorem 3.3. We first consider the sparse case where $k=O(\sqrt{p\log p})$, where we directly apply Theorem 2.3 and Lemma 8.1. We need to show that Assumption 2.2 is satisfied with high probability, and then the results follow from Theorem 2.3. By Lemma 8.1 with $\ell_{\min}=0$, for every $x\in\mathcal{X}_{p,k}$ we pick an arbitrary $y\in\Gamma_{G,v^{*}}(x)$ such that
@@
and set the parent of $x$ to be $y$; this gives a spanning tree with high probability. Note that this automatically establishes the second part of Assumption 2.2 with $\delta=\Omega(\lambda/k^{t})$. Observe that the diameter of the tree is at most $2k$, and the maximum degree of the graph $G_{p,k}$ is at most $kp$. Hence, Assumption 2.2 follows. To upper bound the running time, notice that
@@
for the second part of Assumption 2.2. Thus, by Theorem 2.3, for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-2}}\log p)$ the number of iterations required is at most $O(\beta\lambda k^{2}p)$ with high probability, and for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-3}}\log(p/k))$ it is at most
@@
with high probability. This completes the proof of Theorem 3.3.

### v3

@@ 4.3. The negative result
[Sparse Regression: Negative Result] Assume that for some sufficiently small $c>0$, we have
@@
Then w.h.p. as $k\rightarrow\infty$, for any reversible Markov chain on the graph $G_{p,k}$ with stationary distribution $\pi_{\beta}$, the first time the chain hits $v^{*}$ is $\exp\left(\Omega(k\log(p/k))\right)$.
@@
To prove Theorem 4.5 (see Section 10), we establish a bottleneck for the MCMC methods, by significantly generalizing an Overlap Gap Property result from [26], where the authors assume $n=\Omega(k\log k)$ and $k=p^{o(1)}$. In particular, we generalize to arbitrarily small $n$ and any $k=o(p^{1/3})$.
Remark 4.6.
(Negative result for the posterior temperature) If $\sigma^{2}$ is non-vanishing, then we note that the conditions of Theorem 4.5 are satisfied at the posterior temperature $\beta_{\mathrm{post}}=\frac{1}{2}$ when $n\geq Ck\log(p/k)/(\log(k/\sigma^{2}+1)=\Theta(n_{\mathrm{STATS}})$ for a large enough constant $C$. Therefore Theorem 4.5 implies that any reversible Markov chain on the graph $G_{p,k}$ whose stationary distribution is $\pi_{\mathrm{post}}$ has at least $\exp\left(\Omega(\log(p/k)\right)$ hitting time whenever
$\omega(n_{\mathrm{STATS}})=n=o(n_{\mathrm{ALG}}).$
5. Further Comparison with Previous Work
5.1. Comparison with [2]
For sparse tensor PCA, our results directly compare with [2]. [2] studies the same setting exactly but only under the assumption $t=2$, i.e., in the matrix case, while we study the more general case $t\geq 2$. In this matrix case, they study the same class of low-temperature reversible MCMC methods on the Johnson graph as we consider in this work. They prove that whenever $k=o(p)$ the hitting time of the planted vector is super-polynomial for a part of the conjectured “computationally hard” regime that is when
The authors provide this as evidence for the conjectured computational hardness of sparse (matrix) PCA. Importantly they leave as an open question whether one of these MCMC methods succeed in polynomial-time when other polynomial-time methods are known to work i.e. when $\lambda=\omega(\lambda_{\mathrm{ALG}}(2,k))$.
Our negative results prove that when $k\leq p^{\alpha_{0}}$ for $\alpha_{0}>0$ a small enough, a much wider failure takes place even when $t=2$. Specifically, all considered low-temperature MCMC methods in [2] have super-polynomial hitting time of the planted vector whenever $\lambda=o(k^{3/2}/\sqrt{\log p})$. Moreover, we also provide a positive result saying that the hitting time becomes polynomial for an element of this class (i.e., the Metropolis chain) when $\lambda=\omega(k^{3/2}\sqrt{\log p})$. Hence we pin down the exact threshold, up to logarithmic dependencies, for these MCMC methods to succeed in polynomial-time in sparse PCA, and we highlight that potentially surprisingly it turns out to be much bigger than $\lambda_{\mathrm{ALG}}(2,k)$.
From a technical standpoint, we obtain the mentioned stronger negative MCMC results compared to [2] because of a significantly improved application of the second moment method (see the proof of Proposition 9.4), which is perhaps notable as already [2] used an elaborate second moment method to carry through the same exact argument (when $t=2$). The key new technical ingredient we leverage is some novel “flatness” ideas from random graph theory [3, 28] appropriately adapted to Gaussian tensors. We expect this idea to be broadly useful for similar sparse problems.
5.2. Comparison with [26]
For sparse regression, the most relevant work is [26] that studies the exact same setting as we do. The authors prove that for all $k=o(p)$, if $n=\omega(n_{\mathrm{ALG}})$ a greedy local-search algorithm maximizing the posterior succeeds in polynomial-time. Moreover, they show that if $n=o(n_{\mathrm{ALG}})$ a disconnectivity property which they called Overlap Gap Property for inference takes place, which in particular leads to the failure of the mentioned greedy local search procedure. Importantly, their negative result applies only in the more restricted sparsity regime $k=p^{o(1)}$.
In our work, we prove more general both positive and negative results. For the positive side, we prove that for all $k=o(p)$, a larger family of low-temperature MCMC methods succeeds whenever $n=\omega(n_{\mathrm{ALG}})$. In particular, our positive results include the randomized greedy algorithm maximizing the posterior, as opposed to the vanilla greedy approach of [26]. We highlight that this is a non-trivial extension, which is not always known to be true (see e.g., the relevant discussion in [29] for the planted clique model). On top of that, our positive results hold down to $n=\Theta(n_{\mathrm{ALG}})$ for any $\sigma^{2}=O(k/\log p)$ which is a significant improvement from the required assumption $\sigma^{2}=O(\min\{\frac{\log p}{\log\log p},k\})$ in [26], under a very mild assumption that $k=\omega(\frac{(\log p)^{2}}{\log\log p})$. Finally, we prove that our negative result that for all $k=o(p^{1/3})$ a large family of low-temperature MCMC methods fails when $n=o(n_{\mathrm{ALG}})$, significantly again generalizing the result of [26].
5.3. Comparison with [67] and [73]
In terms of positive results for Bayesian models of sparse regression, there are two other very relevant works on the performance of Markov chains, [67] and [73].
In [67], the authors consider a $k$-sparse regression setting where the goal is to sample from the posterior, given the Zellner hierarchical $g$-prior [72], a deterministic measurement matrix $X$, and i.i.d. Gaussian noise $W$ of variance $\sigma^{2}$. Briefly, instead of considering Markov chains operating only on the exactly $k$-sparse vectors as in our work (which is the parameter space of the problem), they consider a “lifted” state space of all support sets of size below some threshold, and add a regularizer prior to the model that encourages sparsity. They then derive the posterior and study the Metropolis chain [^4] The graph in this case has two states connected if their Hamming distance is at most two, rather than exactly two as in our case. to sample from it. They go on to derive polynomial mixing time bounds and hitting time guarantees of the planted vector of the Metropolis chain sampling from the posterior. Their result hold under certain structural deterministic assumptions on the feature matrix $X$.
Comparing with our results where $X$ has i.i.d. standard Gaussian entries, we show that these structural conditions on $X$ are not satisfied with high probability when $n=o(k\sigma)$, where $\sigma^{2}$ is the noise variance (recall $n_{\mathrm{ALG}}=\Theta(k\log(p/k))$. We present a proof of this claim in Section A.3. Note that when $\sigma^{2}=\Theta(k/\log p)$ (the maximum noise level allowed for our positive result to hold), this implies that the required sample size for the MCMC algorithm in [67] to succeed is $n=\Omega\left(k^{1.5}/\sqrt{\log p}\right)$, which significantly larger than the conjectures algorithmic threshold $n_{\mathrm{ALG}}=\Theta(k\log(p/k))$ which is achieved by our analysis of the Metropolis process using our main positive result Theorem 2.3.
The same setting of sparse regression with deterministic $X$ and Gaussian noise $W$ as in [67] is considered in [73]. Here the authors study a different Metropolis chain but on the same state space as [67] (that is of supports of arbitrary size below some threshold), and prove that, under assumptions, it mixes with the posterior rapidly. Of note, the final mixing time bound obtained is independent of $p$. A key feature of their Metropolis chain is the use of an informed proposal method, where subsets of variables that have a higher posterior probability are more likely to be proposed as transitions. Such informed proposal schemes have been studied in several previous works [71, 70, 30]. This contrasts with the usual version of Metropolis, where proposed transitions are chosen in an uninformed manner, such as uniformly (the approach taken both in our work and in [67]). However, informed proposals carry a significantly higher computational cost than uninformed ones, as they generally require calculating the full posterior probability of each neighbor of the current state of the chain. This computational cost can be up to $O(pk^{2})$ at each iteration, rather than $O(nk)$ in our case (the cost of evaluating $H$ at a single point). The total computational complexity of the algorithm in [73] therefore has a linear dependence on $p$, identical to our result.
We also note that the assumptions on $X$ made in [73], while deterministic assumptions, are slightly weaker than those in [67]. Rather than making the same structural assumptions on the design matrix $X$, the authors show that if there exists a path (in Hamming distance) from any set of coordinates (of arbitrary sparsity) to the true set, such that the posterior probability is suitably increasing at each step, then the chain mixes quickly. Then they prove that the structural assumptions in [67] indeed imply their required condition. We highlight that this assumption is in fact very similar to the general assumption we propose for arbitrary state space 2.2.
We highlight, however, some key further differences between our results and that in [73]. First, as mentioned above we analyze the uninformed Metropolis process on the $k$-sparse vectors, while they do anayze the informed Metropolis process on sparse vectors of arbitrary size below some threshold. Second, the authors do not prove that their “ascending paths” assumptions on this extended state space is satisfied even in some relatively simple specific examples such as the case of i.i.d. Gaussian $X$, instead they show that their assumptions are implied by the structural assumptions in [67]. As argued above, these structural assumptions of [67] can fail to hold for i.i.d. Gaussian data when $n=\Theta(n_{\mathrm{ALG}})$ and $\sigma^{2}=\Theta(k/\log p)$. On the other hand, we do establish that our 2.2 does indeed hold down to $n=\Theta(n_{\mathrm{ALG}})$ when we focus on the $k$-sparse state space.
@@ 6.2.1. Specification
Our implementation here uses $p=20000$, $k=18$, $\sigma^{2}=0.05$, and $\beta=100$. We then simulate the Metropolis chain for varying values of the sample size $n$. The information-theoretic threshold here is given by [26]
$n_{\mathrm{STATS}}=\frac{2k\log(p/k)}{\log\left(\frac{k}{\sigma^{2}}+1\right)}\approx 43$
Our positive result in Theorem 4.2 provides the correct order for polynomial time convergence, but not the exact constant. We therefore compare to the algorithmic threshold for LASSO, where the exact threshold for success is given by [64]
$n_{\mathrm{LASSO}}=(2k+\sigma^{2})\log(p/k)\approx 253$
@@ 6.2.2. Results
The results of these simulations are displayed in Fig. 4. We again observe a clear phase transition when the sample size $n$ crosses the algorithmic threshold. When $n<n_{\mathrm{LASSO}}$, most samples time out after a million iterations having failed to find $v^{*}$. Above $n_{\mathrm{LASSO}}$, the chain quickly finds $v^{*}$. We can see these dynamics in more detail in Fig. 5. Here we group simulations by their value of $n$. Then at each time, we plot the median proportion of $v^{*}$ recovered at that time within that group. We observe that for values of $n$ above 300, we achieve very rapid convergence. Conversely, below 175 the chain does not appear to be converging at all, despite being significantly above the information-theoretic threshold $n_{\mathrm{STATS}}\approx 43$.
We recognize that the correct constant for the MCMC algorithmic threshold may not be the same as for LASSO, and when $n$ is relatively small as it is here, that can affect the results. Nevertheless, we do see a rapid speed up in convergence when the LASSO threshold is surpassed.
@@ 7. Proof of General Positive Result
For part (2), we use the standard canonical paths argument for mixing (for example, see chapter 5 of [37]). For any $x,y\in\mathcal{X}$ we define $\gamma_{x,y}$ to be the unique path connecting them in $T$. Note that $|\gamma_{x,y}|\leq D$. Now for each edge $(a,b)$ in $T$, we upper bound the conductance $t(a,b)$ of it, defined as
@@
Thus, the conductance of the Metropolis chain on $G$ is at most $O(D\Delta)$ and therefore the mixing time is at most $O(D\Delta\left(\log|\mathcal{X}|+\beta\mathcal{R}_{H}\right))$ using standard results, see [18, 60] or [37, Theorem 5.2]. Now, given the mixing time upper bound, part (2) follows directly from part (1).
Lastly, we consider the case where $\beta\geq\frac{2}{\delta}(\log|\mathcal{X}|+\log(\frac{40\mathcal{R}_{H}}{\delta}))$ and prove part (3). We argue that in such a low-temperature regime, the behavior of the Metropolis chain is very similar as the Randomized Greedy algorithm with high probability.
We call a step of the Metropolis chain to be “good” if, at current state $X_{t-1}=x\neq v^{*}$, it selects $y\in\Gamma_{G}(x)$ which is the parent of $x$ in $T$; note that since $H(y)>H(x)$ by 2.2, the chain will then move to $y$, i.e., $X_{t}=y$. For completeness, if the current state is $X_{t-1}=v^{*}$ then we always call this step good no matter which next state it picks or whether it moves. Thus, each step has a probability at least $1/\Delta$ to be good. A simple application of the Chernoff bounds yields that, if we run the Metropolis chain for $T=\omega(\frac{\Delta\mathcal{R}_{H}}{\delta})$ steps then – since for large enough $n$ it must hold $T\geq\frac{20\Delta\mathcal{R}_{H}}{\delta}$ – we have that at least $\frac{4\mathcal{R}_{H}}{\delta}$ of the first $T$ steps of the Metropolis chain are good with probability at least $0.9$.
Let $m^{\prime}$ be the number of good steps. Suppose the first $m=\min\left\{\frac{4\mathcal{R}_{H}}{\delta},m^{\prime}\right\}$ good steps happen at time $1\leq t_{1}<t_{2}<\cdots<t_{m}\leq T$; we also set $t_{0}=0$. Since the transition matrix, conditional on no move between time $t_{i-1}$ and $t_{i}$ is good, is still time-reversible, it holds that
@@
Now, by the union bound, with probability at least $0.8$, we have that (1) there are at least $m^{\prime}\geq\frac{4\mathcal{R}_{H}}{\delta}$ good steps; (2) for all $1\leq i\leq m=\frac{4\mathcal{R}_{H}}{\delta}$ it holds $H(X_{t_{i}-1})>H(X_{t_{i-1}})-\frac{\delta}{2}$. We claim that under (1) and (2), it must hold that $X_{t_{i}-1}=v^{*}$ for some $1\leq i\leq m$, and the theorem then follows. Suppose for sake of contradiction that $X_{t_{i}-1}\neq v^{*}$ for all $i$. Then, by 2.2 we have $H(X_{t_{i}})\geq H(X_{t_{i}-1})+\delta>H(X_{t_{i-1}})+\frac{\delta}{2}$. It follows that $H(X_{t_{m}})>H(X_{t_{0}})+\frac{\delta m}{2}=H(X_{0})+2\mathcal{R}_{H}$, contradicting $H(X_{t_{m}})-H(X_{0})\leq\mathcal{R}_{H}$ by definition. ∎
@@ 8.1. Proof of Positive Results for the Gaussian Additive Model
Here we apply Theorem 2.3 to prove Theorem 3.3 under 3.2. The first step is to show that 3.2 implies 2.2.
8.1.1. Verifying 2.2.
@@
In the following lemma, we verify 2.2 under 3.2.
@@
Assume that 3.2 holds with constants $\varepsilon\in(0,1)$ and $t\geq 1$. For all positive integers $p,k,\ell_{\min}\in\mathbb{N}^{+}$ such that $2/\varepsilon\leq k\leq p$ and $\ell_{\min}\leq\varepsilon k$, if
@@
We deduce from 3.2 that
@@
We want to show that there exists at least half of $y\in\Gamma_{G,v^{*}}(x)$ that satisfies $|Z^{x}_{y}|=O\left(\sqrt{(\log p)/k}\right)$.
Observe that each $Z^{x}_{y}$ is a Gaussian with mean zero and variance given by
@@
where we use the convexity of $f$ and item (b) of 3.2. Thus, we have
@@
where we apply the convexity of $f$ and item (c) of 3.2. Meanwhile, the number of $y^{\prime}\in\Gamma_{G,v^{*}}(x)$ such that $\left\langle y,y^{\prime}\right\rangle=k-1$ is $2(k-\ell-1)$, and for such $w$ it holds
@@
by the convexity of $f$ and item (b) of 3.2. Finally, if $y^{\prime}=y$ then $\left\langle y,y^{\prime}\right\rangle=k$ and
@@
Lemma 8.2 (Hanson–Wright Inequality [57, Theorem 1.1]).
@@
We are now ready to prove Theorem 3.3. We first consider the sparse case where $k=O(\sqrt{p\log p})$, where we directly apply Theorem 2.3 and Lemma 8.1. We need to show that 2.2 is satisfied with high probability, and then the results follow from Theorem 2.3. By Lemma 8.1 with $\ell_{\min}=0$, for every $x\in\mathcal{X}_{p,k}$ we pick an arbitrary $y\in\Gamma_{G,v^{*}}(x)$ such that
@@
and set the parent of $x$ to be $y$; this gives a spanning tree with high probability. Note that this automatically establishes the second part of 2.2 with $\delta=\Omega(\lambda/k^{t})$. Observe that the diameter of the tree is at most $2k$, and the maximum degree of the graph $G_{p,k}$ is at most $kp$. Hence, 2.2 follows.
To upper bound the running time, notice that
@@
For the final statement of Theorem 3.3, we note that by Lemma 8.1, for any neighbor $y$ of $v^{*}$, it holds with high probability that $H(v^{*})-H(y)=\Omega\left(\frac{\lambda}{k^{t}}\right)$. Therefore, once the Randomized Greedy algorithm (i.e. when $\beta=\infty$) hits $v^{*}$, it remains there for all future iterations with high probability.
@@
for the second part of 2.2. Thus, by Theorem 2.3, for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-2}}\log p)$ the number of iterations required is at most $O(\beta\lambda k^{2}p)$
with high probability, and for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-3}}\log(p/k))$ it is at most
@@
with high probability. Again, for the final statement of Theorem 3.3, we note that for any neighbor $y$ of $v^{*}$, it holds with high probability that $H(v^{*})-H(y)=\Omega\left(\frac{\lambda k^{t-2}}{p^{t-1}}\right)$. Therefore, once the Randomized Greedy algorithm (i.e. when $\beta=\infty$) hits $v^{*}$, it remains there for all future iterations with high probability. This completes the proof of Theorem 3.3.